OpenStudy (anonymous):

: Find the equation of the line below. Arrange your answer in the form y = mx + b, where b is the constant. where the coordinates are (1,2) and (-1,-2)

OpenStudy (isaiah.feynman):

Where's the line?

OpenStudy (anonymous):

the coordinates are (2,1) and (-2,-1)

OpenStudy (anonymous):

y=2x+0, or just y=2x. hope this helped

OpenStudy (anonymous):

thanks what about this one? Find the equation of the line below. Arrange your answer in the form y = mx + b, where b is the constant. where the coordinates are (-2,2) and (2,-2)

OpenStudy (isaiah.feynman):

@magan the question is the same format as the first, that should give you a clue as to go about the problem!

OpenStudy (anonymous):

http://www.coolmath.com/algebra/08-lines/06-finding-slope-line-given-two-points-01.htm. a little childish, but it helps y=-1x+0 or y= -x